# PepTune: De Novo Generation of Therapeutic Peptides with Multi-Objective-Guided Discrete Diffusion

### v1

@@ 4.5 Evaluation
Peptide Validity Filter. Among the sequential representations of peptides including amino acid sequences, HELM [104], and SMILES [32], SMILES is the most intricate representation of peptide sequences. Although this enables the representation of non-natural amino acids, diverse side-chain, and backbone modifications, and cyclic peptides, it also means that the vast majority of SMILES strings are not synthesizable peptides. Therefore, we devised an algorithm that determines whether a SMILES string is a valid peptide, characterized by peptide bonds and central carbon atoms. The filter first checks if the SMILES sequence is a valid molecule using RDKit [105]. Then we check whether the generated SMILES is peptide or not by its peptide bonds and N-methylated peptide bonds and determine whether the peptides are cyclic.
@@
We use regular expressions to detect bond patterns for peptide bonds, N-methylated peptide bonds, reversed peptide bonds, and ester bonds, along the sequence to split the sequence into several segments with a bond before and after each segment. The filter then checks each segment for chemical modifications based on their bond type, including N-methylation (N-Me) and O-linked glycosylation. The remaining segment is matched to the corresponding amino acid side chains for either natural or non-natural sources. The tool is freely available on HuggingFace: https://huggingface.co/spaces/ChatterjeeLab/SMILES2PEPTIDE.
@@
Generation Quality Metrics. To evaluate the generation quality of our unconditional MDLM, PepMDLM, and our MCTS-guided MDLM, PepTune, we leverage the Moses metrics [36]. Validity is the percentage of the generated SMILES resulting in valid peptide SMILES based on our validity filter that first checks if the string is a valid molecule SMILES, then checks for the presence of unmodified and modified peptide bonds within the sequence. Uniqueness is the fraction of distinct peptide SMILES sequences among the valid peptide SMILES. Diversity is calculated as one minus the Tanimoto similarity between the Morgan fingerprints for every pair of generated sequences. Similarity to nearest neighbor (SNN) takes the maximum Tanimoto similarity score for each generated sequence with any other sequence and averages across all generated sequences to measure the similarity on average that a sequence to its closest neighbor.
@@
Due to the limit of memory and CPU time required to load all the training dataset of 11 million peptide SMILES, we chose to sample a subset of 1000 batches randomly ($\sim$100k sequences) for diversity and SNN calculation. To assess the novelty of generated sequences, we employed Shannon entropy [106] to quantify the SMILES token randomness between 100 conditionally and 100 unconditionally generated sequences and the same randomly sampled 1000 subsets from the training set. Then Kullback-Leibler (KL) divergence was also used to evaluate divergence across token distributions. The equations for all metrics are provided in Appendix C.5.
@@
Docking. For generated SMILES that were identified as peptides by our filter and fulfilled the targeted properties, we used Autodock Vina [48] (v 1.1.2) for in silico docking experiment between the target proteins and the ligands. Targets were preprocessed with MGITools [107] (v 1.5.7) and the conformations of the SMILES were optimized by ETKDG from RDKit [48, 108]. The final results were visualized in PyMol [109] (v 3.1) with polar contacts within 3.5 Å being labeled.
@@ 5 Declarations
Acknowledgments. We thank the Duke Compute Cluster, Pratt School of Engineering IT department, and Mark III Systems, for providing database and hardware support that have contributed to the research reported within this manuscript. We thank Alexander Tong for reviewing the theoretical formulations of PepTune. We also thank Sophia Vincoff and Lauren Hong for assistance with figure generation.
@@ A.1 Continuous-Time Discrete Diffusion
We denote a token in a sequence from the dataset as a one-hot vector $\mathbf{x}_{0}^{(\ell)}\in\{0,1\}^{|\mathcal{V}|}$. The discrete-time forward diffusion process involves applying discrete categorical noise to $\mathbf{x}_{0}$ at varying degrees based on a noise schedule $\sigma(t)$ for a total of $T$ time steps ranging from no noise at $t=0$ to maximum noise at $t=1$. To clearly distinguish each step, we denote the $n$th transition in the forward pass as the transition from $s(n)$ to $t(n)$, where $s(n)=\frac{n-1}{T}$ and $t(n)=\frac{n}{T}$. The marginal noise that transforms the sequence $\mathbf{z}_{s(n)}$ at time $s(n)$ to a progressively noisier sequence $\mathbf{z}_{t(n)}$ the next time step $t(n)=s(n)+\frac{1}{T}$ is given by a $|\mathcal{V}|\times|\mathcal{V}|$ transition matrix $\mathbf{Q}_{t|s}=\sigma(t)\mathbf{I}_{|\mathcal{V}|}+(1-\sigma(t))\mathbf{1}\mathbf{\pi}^{\top}$, where the $(i,j)$th entry denotes the probability of transitioning from token $i$ to token $j$ at each position in the sequence.
@@
For the continuous-time forward pass diffusion process where $T\to\infty$ and $\frac{1}{T}\to 0$, we can take the limit as $T\to\infty$ to derive an expression for the continous-time transition probability, $\alpha_{t}$.
@@
Following Austin et al. [19] and substituting the marginal and cumulative probability distributions, we derive the true reverse distribution from $t\to s$ conditioned on a clean sequence $\mathbf{x}_{0}$ as
@@
$=\text{Cat}\left(\mathbf{z}_{s};\frac{[\alpha_{t|s}\mathbf{z}_{t}+(1-\alpha_{t|s})\mathbf{1}\mathbf{\pi}^{\top}\mathbf{z}]\odot[\alpha_{s}\mathbf{z}_{t}+(1-\alpha_{s})\mathbf{z}_{s}]}{\mathbf{z}_{t}^{\top}\mathbf{Q}_{t}^{\top}\mathbf{x}_{0}}\right)$ (40)
The numerator of the discrete reverse posterior is the element-wise product of $|\mathcal{V}|$-dimensional vectors representing the marginal probability distribution of sampling $\mathbf{z}_{t}$ given $\mathbf{z}_{s}$ and the total probability distribution of sampling $\mathbf{z}_{s}$ from the original clean sequence $\mathbf{x}_{0}$. The denominator is a scalar probability of the specific token $\mathbf{z}_{t}$ being drawn from the noisy probability distribution at time $t$.
@@ A.2 Continuous-Time Negative Evidence Lower Bound (NELBO)
The objective of denoising diffusion probabilistic models (DDPMs) [64] is to iteratively sample slightly less noisy intermediate sequences $\mathbf{z}_{t}$ until obtaining a clean sequence $\mathbf{x}$ that has a high probability of being drawn from the distribution $p(\mathbf{x}_{0})$ of all possible valid sequences. Since $p(\mathbf{x})$ is unknown, we compute an Evidence Lower Bound (ELBO) of $\log p(\mathbf{x}_{0})$ such that by maximizing the ELBO, we are maximizing $\log p(\mathbf{x}_{0})$ that measures how well the model can generate true samples $\mathbf{x}_{0}$ from its learned distribution $p_{\theta}(\mathbf{x_{0}|\mathbf{z}_{t}})$. To convert this into a loss minimization objective, we define the loss function as the Negative ELBO (NELBO). First, we can compute $p(\mathbf{x}_{0})$ by integrating over the joint probability of all possible paths of intermediate states from the noisy state $\mathbf{x}_{T}$ at $t=T$ to the clean sample $x_{0}$ at $t=0$.
@@
$=\log\left(\mathbb{E}_{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\left[\frac{p(\mathbf{x}_{0:T})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right]\right)$ (41)
Using Jenson’s inequality, we move the logarithm inside the expectation to get the ELBO for $\log p(\mathbf{x}_{0})$. Further, we split the terms associated with the forward noising process $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$ and the reverse denoising model $p(\mathbf{x}_{0:T})$ into reconstruction term, the intermediate diffusion terms, and the prior term.
$\mathcal{L}_{\text{NELBO}}$ $=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log\frac{p(\mathbf{z}_{0:T})}{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\right]$
$=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log\frac{p(\mathbf{x}_{0}|\mathbf{z}_{t(1)})p(\mathbf{z}_{t(T)})\prod_{n=1}^{T-1}p(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})}{q(\mathbf{x}_{T}|\mathbf{x}_{T-1})\prod_{n=1}^{T-1}q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})}\right]$
$=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log p(\mathbf{x}_{0}|\mathbf{x}_{1})-\log\frac{p(\mathbf{x}_{T})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})}-\log\frac{\prod_{t=1}^{T-1}p(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})}{\prod_{n=1}^{T-1}q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})}\right]$
$=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log p(\mathbf{x}_{0}|\mathbf{z}_{t(1)})-\log\frac{p(\mathbf{z}_{t(T)})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})}-\sum_{n=1}^{T-1}\log\frac{p(\mathbf{z}_{s(n)}|\mathbf{z}_{s(n)})}{q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})}\right]$
$=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\bigg{[}-\log p(\mathbf{x}_{0}|\mathbf{z}_{t(1)})\bigg{]}+\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log\frac{p(\mathbf{z}_{t(T)})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})}\right]$
$\quad+\sum_{n=1}^{T-1}\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\bigg{[}-\log\frac{p(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})}{q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})}\bigg{]}$
$=\underbrace{\mathbb{E}_{q(\mathbf{z}_{t(1)}|\mathbf{x}_{0})}\bigg{[}-\log p(\mathbf{x}_{0}|\mathbf{z}_{t(1)})\bigg{]}}_{\text{reconstruction loss}}+\underbrace{\mathbb{E}_{q(\mathbf{z}_{t(T)},\mathbf{z}_{s(T)}|\mathbf{x}_{0})}\left[-\log\frac{p(\mathbf{z}_{t(T)})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})}\right]}_{\text{prior loss}}$
@@
Now, we can take the limit for each of the loss terms as $T\to\infty$ to derive the general form of the continuous-time MDLM objective.
@@
$p(\mathbf{z}_{t(1)}|\mathbf{x}_{0})=\text{Cat}(\mathbf{z}_{t(1)};\vec{\alpha}^{\top}_{t(1)}\mathbf{x}_{0}\mathbf{x}_{0}+(1-\vec{\alpha}^{\top}_{t(1)}\mathbf{x}_{0})\mathbf{m})$ (43)
@@
$\alpha_{t}(\mathbf{x}_{0})\mathbf{x}_{0}+(1-\alpha_{t}(\mathbf{x}_{0}))\mathbf{m}$ =$=\begin{cases}\left(1-\frac{1}{T^{w}}\right)\mathbf{x}_{0}+\frac{1}{T^{w}}\mathbf{m}&\mathbf{x}_{0}=\mathbf{b}\\ \left(1-\frac{1}{T}\right)\mathbf{x}_{0}+\frac{1}{T}\mathbf{m}&\mathbf{x}_{0}\neq\mathbf{b}\end{cases}$ (44)
@@
Prior Loss $\mathcal{L}_{\text{prior}}$. The prior loss measures the first reverse transition from the fully masked sequence $\mathbf{z}_{t(T)}$ to a slightly unmasked sequence $\mathbf{z}_{t(T-1)}$.
@@
Since $t(T)=1$ and $t(T-1)=\frac{T-1}{T}$, we can derive
$q(\mathbf{z}_{t(T)}|\mathbf{x}_{0})$ $=\text{Cat}\big{(}\mathbf{z}_{t(T)}\;;\;\alpha_{t}(\mathbf{x}_{0})\mathbf{x}_{0}+(\mathbf{1}-\alpha_{t}(\mathbf{x}_{0}))\mathbf{m}\big{)}$
@@
It follows that the prior distribution $p_{\theta}(\mathbf{z}_{t(T)})=\text{Cat}\big{(}\mathbf{z}_{t(T)};\mathbf{m}\big{)}$ and the KL divergence reduces to 0.
Diffusion Loss $\mathcal{L}_{T}$. The diffusion loss measures the error of each reverse diffusion step for $T$ time steps. That is, it measures the consistency of the learned reverse transitions with the forward pass diffusion.
@@
Since the objective is to accurate predict $\mathbf{z}_{s(n)}$ given $\mathbf{z}_{t(n)}$, we cannot condition on the term $\mathbf{z}_{s}$. Instead, we can condition $q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})$ on $\mathbf{x}_{0}$ and use Baye’s theorem to derive

### v2

@@ 4.5 Evaluation
Peptide Validity Filter. Among the sequential representations of peptides including amino acid sequences, HELM [106], and SMILES [32], SMILES is the most intricate representation of peptide sequences. Although this enables the representation of non-natural amino acids, diverse side-chain, and backbone modifications, and cyclic peptides, it also means that the vast majority of SMILES strings are not synthesizable peptides. Therefore, we devised an algorithm that determines whether a SMILES string is a valid peptide, characterized by peptide bonds and central carbon atoms. The filter first checks if the SMILES sequence is a valid molecule using RDKit [107].
@@
Then, we use regular expressions to detect bond patterns for peptide bonds, N-methylated peptide bonds, reversed peptide bonds, and ester bonds, along the sequence to split the sequence into several segments with a bond before and after each segment. The filter checks each segment for chemical modifications based on their bond type, including N-methylation (N-Me) and O-linked glycosylation. The remaining segment is matched to the corresponding natural or non-natural amino acid side chains (Algorithm 8). Our filter is capable of detecting a library of over 200 nAAs from SwissSidechain [40] and can classify a peptide SMILES as cyclic or non-cyclic. The tool is freely available on HuggingFace: https://huggingface.co/spaces/ChatterjeeLab/SMILES2PEPTIDE.
@@
Generation Quality Metrics. To evaluate the generation quality of our unconditional MDLM, PepMDLM, and our MCTS-guided MDLM, PepTune, we leverage the Moses metrics, including validity, uniqueness, diversity, similarity to nearest neighbor (SNN), and novelty [36]. Validity is the percentage of generated sequences that are valid peptide SMILES based on our peptide validity filter. Uniqueness is the fraction of distinct peptide SMILES sequences among the valid peptide SMILES. Diversity is calculated as one minus the average Tanimoto similarity between the Morgan fingerprints for every pair of generated sequences. Similarity to nearest neighbor (SNN) takes the maximum Tanimoto similarity score for each generated sequence with the sequences in the training dataset and averages across all generated sequences to measure the average similarity of a generated sequence with its most similar neighbor.
@@
Due to the limit of memory and CPU time required to load all the training dataset of 11 million peptide SMILES, we chose to sample a subset of 1000 batches randomly ($\sim$100k sequences) for novelty and SNN calculation. To assess the novelty of generated sequences, we employed Shannon entropy [108] to quantify the SMILES token randomness between 100 PepTune-generated and 100 PepMDLM-generated sequences and the same randomly sampled 1000 subsets from the training set. Then Kullback-Leibler (KL) divergence was also used to evaluate divergence across token distributions from the generated peptides compared to the training data. The equations for all metrics are provided in Appendix C.5.
@@
Docking. For valid generated peptide SMILES with non-dominated scores across objectives, we used Autodock Vina [48] (v 1.1.2) for in silico docking of the peptide binders to their target proteins (Appendix 9) to confirm binding affinity. Targets were preprocessed with MGITools [109] (v 1.5.7) and the conformations of the SMILES were optimized by ETKDG from RDKit [48, 110]. The final results were visualized in PyMol [111] (v 3.1), where the residues in the protein targets with polar contacts to the peptide binder with distances closer than 3.5 Å are annotated.
@@ 5 Declarations
Acknowledgments. We thank the Duke Compute Cluster, Pratt School of Engineering IT department, and Mark III Systems, for providing database and hardware support that has contributed to the research reported within this manuscript. We thank Alexander Tong for reviewing the theoretical formulations of PepTune. We also thank Sophia Vincoff and Lauren Hong for their assistance with figure generation.
@@ A.1 Continuous-Time Discrete Diffusion
We denote a token in a sequence from the dataset as a one-hot vector $\mathbf{x}_{0}^{(\ell)}\in\{0,1\}^{|\mathcal{V}|}$. The discrete-time forward diffusion process involves applying categorical noise to $\mathbf{x}_{0}$ at varying degrees based on a noise schedule $\sigma(t)$ for a total of $T$ time steps ranging from no noise at $t=0$ to maximum noise at $t=1$. To clearly distinguish each step, we denote the $n$th transition in the forward pass as the transition from $s(n)$ to $t(n)$, where $s(n)=\frac{n-1}{T}$ and $t(n)=\frac{n}{T}$. The marginal noise that transforms the sequence $\mathbf{z}_{s(n)}$ at time $s(n)$ to a progressively noisier sequence $\mathbf{z}_{t(n)}$ at the next time step $t(n)=s(n)+\frac{1}{T}$ is given by a $|\mathcal{V}|\times|\mathcal{V}|$ marginal transition matrix $\mathbf{Q}_{t|s}=\sigma(t)\mathbf{I}_{|\mathcal{V}|}+(1-\sigma(t))\mathbf{1}\mathbf{\pi}^{\top}$, where the $(i,j)$th entry denotes the probability of transitioning from token $i$ to token $j$ at each position in the sequence.
@@
For the continuous-time forward pass diffusion process where $T\to\infty$ and $\frac{1}{T}\to 0$, we can take the limit as $T\to\infty$ to derive an expression for the continuous-time transition probability, $\alpha_{t}$.
@@
Following Austin et al. [19] and substituting the marginal and cumulative probability distributions, we derive the true reverse transition from $t\to s$ conditioned on a clean sequence $\mathbf{x}_{0}$ as
@@
$=\text{Cat}\left(\mathbf{z}_{s};\frac{[\alpha_{t|s}\mathbf{z}_{t}+(1-\alpha_{t|s})\mathbf{1}\mathbf{\pi}^{\top}\mathbf{z}_{t}]\odot[\alpha_{s}\mathbf{x}_{0}+(1-\alpha_{s})\mathbf{\pi}]}{\alpha_{t}\mathbf{z}_{t}^{\top}\mathbf{x}_{0}+(1-\alpha_{t})\mathbf{z}_{t}^{\top}\mathbf{\pi}}\right)$ (37)
where the numerator is the element-wise product of $|\mathcal{V}|$-dimensional vectors representing the marginal probability distribution of sampling $\mathbf{z}_{t}$ given $\mathbf{z}_{s}$ and the cumulative probability distribution for $\mathbf{z}_{s}$ from the original clean sequence $\mathbf{x}_{0}$. The denominator is a scalar probability of the specific token $\mathbf{z}_{t}$ being drawn from the noisy probability distribution at time $t$.
@@ A.2 Continuous-Time Negative Evidence Lower Bound (NELBO)
The objective of denoising diffusion probabilistic models (DDPMs) [64] is to iteratively sample slightly less noisy intermediate sequences $\mathbf{z}_{t}$ until obtaining a clean sequence $\mathbf{x}$ that has a high probability of being drawn from the data distribution $p(\mathbf{x}_{0})$. To train a model that accurately samples from $p(\mathbf{x}_{0})$, we maximize the Evidence Lower Bound (ELBO) of $\log p_{\theta}(\mathbf{x}_{0})$ which measures how accurately the model parameterized by $\theta$ generates true samples $\mathbf{x}_{0}$ given a corrupted sequence $\mathbf{z}_{T}$ by iterative sampling from the reverse posterior $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})$. The ELBO is maximized when $p_{\theta}(\mathbf{x}_{0})=1$ and $\log(p_{\theta}(\mathbf{x}_{0}))=0$ for all sequences $\mathbf{x}_{0}$ in the dataset, which supports the objective of accurately generating sequences from the data distribution. To convert this into a loss minimization objective, we define the loss function as the negative ELBO (NELBO). First, we compute $\log p_{\theta}(\mathbf{x}_{0})$ by integrating over the joint probability of all possible paths of intermediate states from the noisy state $\mathbf{z}_{T}$ at $t=T$ to the clean sample $x_{0}$ at $t=0$, denoted by $p_{\theta}(\mathbf{x}_{0:T})$. Since our goal is to reverse the forward masking of the clean sample $x_{0}$ from all time steps, we introduce an encoder term $q(\mathbf{z}_{1:T}|\mathbf{x}_{0})$ denoting the combined distribution of obtaining any noisy sequence between times $t=1$ to $t=T$ from the clean sequence $\mathbf{x}_{0}$.
@@
$=\log\left(\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[\frac{p_{\theta}(\mathbf{z}_{0:T})}{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\right]\right)$ (38)
where $\mathbf{z}_{0:T}$ includes $\mathbf{x}_{0}$.
Using Jenson’s inequality, we move the logarithm inside the expectation and reverse the sign to get the NELBO for $\log p_{\theta}(\mathbf{x}_{0})$. We split the terms associated with the forward noising process $q(\mathbf{z}_{1:T}|\mathbf{x}_{0})$ and the reverse denoising model $p_{\theta}(\mathbf{z}_{0:T})$ into reconstruction term, the prior term, and the intermediate diffusion term.
$\mathcal{L}_{\text{NELBO}}$ $=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log\frac{p_{\theta}(\mathbf{z}_{0:T})}{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\right]$
$=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log\frac{p_{\theta}(\mathbf{x}_{0}|\mathbf{z}_{t(1)})p_{\theta}(\mathbf{z}_{t(T)})\prod_{n=1}^{T-1}p_{\theta}(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{t(T-1)})\prod_{n=1}^{T-1}q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})}\right]$
$=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{z}_{t(1)})-\log\frac{p_{\theta}(\mathbf{z}_{t(T)})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})}-\log\frac{\prod_{t=1}^{T-1}p_{\theta}(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})}{\prod_{n=1}^{T-1}q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})}\right]$
$=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{z}_{t(1)})-\log\frac{p_{\theta}(\mathbf{z}_{t(T)})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})}-\sum_{n=1}^{T-1}\log\frac{p_{\theta}(\mathbf{z}_{s(n)}|\mathbf{z}_{s(n)})}{q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})}\right]$
$=\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\bigg{[}-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{z}_{t(1)})\bigg{]}+\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\left[-\log\frac{p_{\theta}(\mathbf{z}_{t(T)})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})}\right]$
$\quad+\sum_{n=1}^{T-1}\mathbb{E}_{q(\mathbf{z}_{1:T}|\mathbf{x}_{0})}\bigg{[}-\log\frac{p_{\theta}(\mathbf{z}_{s(n)}|\mathbf{z}_{t(n)})}{q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})}\bigg{]}$
$=\underbrace{\mathbb{E}_{q(\mathbf{z}_{t(1)}|\mathbf{x}_{0})}\bigg{[}-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{z}_{t(1)})\bigg{]}}_{\text{reconstruction loss}}+\underbrace{\mathbb{E}_{q(\mathbf{z}_{t(T)},\mathbf{z}_{s(T)}|\mathbf{x}_{0})}\left[-\log\frac{p_{\theta}(\mathbf{z}_{t(T)})}{q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})}\right]}_{\text{prior loss}}$
@@
Now, we can take the limit for each of the loss terms as $T\to\infty$ to derive the continuous-time MDLM objective.
@@
$p(\mathbf{z}_{t(1)}|\mathbf{x}_{0})=\text{Cat}(\mathbf{z}_{t(1)};\alpha_{t(1)}(\mathbf{x}_{0})\mathbf{x}_{0}+(1-\alpha_{t(1)}(\mathbf{x}_{0}))\mathbf{m})$ (40)
@@
$\alpha_{t(1)}(\mathbf{x}_{0})\mathbf{x}_{0}+(1-\alpha_{t(1)}(\mathbf{x}_{0}))\mathbf{m}$ =$=\begin{cases}\left(1-\frac{1}{T^{w}}\right)\mathbf{x}_{0}+\frac{1}{T^{w}}\mathbf{m}&\mathbf{x}_{0}=\mathbf{b}\\ \left(1-\frac{1}{T}\right)\mathbf{x}_{0}+\frac{1}{T}\mathbf{m}&\mathbf{x}_{0}\neq\mathbf{b}\end{cases}$ (41)
@@
Prior Loss $\mathcal{L}_{\text{prior}}$. The prior loss measures the first reverse transition from the fully masked sequence $\mathbf{z}_{t(T)}$ to a slightly unmasked sequence $\mathbf{z}_{s(T)}$.
@@
Since $t(T)=1$, we have $\alpha_{t(T)}(\mathbf{x}_{0})=1-1=0$. Therefore, we derive
$q(\mathbf{z}_{t(T)}|\mathbf{x}_{0})$ $=\text{Cat}\big{(}\mathbf{z}_{t(T)}\;;\;\alpha_{t(T)}(\mathbf{x}_{0})\mathbf{x}_{0}+(1-\alpha_{t(T)}(\mathbf{x}_{0}))\mathbf{m}\big{)}$
$=\text{Cat}\big{(}\mathbf{z}_{t(T)}\;;\;0\mathbf{x}_{0}+(1-0)\mathbf{m}\big{)}$
@@
Since all sequences are completely masked at time $T$, it follows that the marginal distribution $q(\mathbf{z}_{t(T)}|\mathbf{z}_{s(T)})=\text{Cat}\big{(}\mathbf{z}_{t(T)};\mathbf{m}\big{)}$ and the prior distribution $p_{\theta}(\mathbf{z}_{t(T)})=\text{Cat}\big{(}\mathbf{z}_{t(T)};\mathbf{m}\big{)}$, so the KL divergence reduces to 0.
Diffusion Loss $\mathcal{L}_{T}$. The diffusion loss measures the consistency of each predicted reverse transition with the forward marginal transition for all $T$ time steps.
@@
Since the objective is to accurate predict $\mathbf{z}_{s(n)}$ given $\mathbf{z}_{t(n)}$, we cannot condition on the term $\mathbf{z}_{s(n)}$. Instead, we can condition $q(\mathbf{z}_{t(n)}|\mathbf{z}_{s(n)})$ on $\mathbf{x}_{0}$ and use Baye’s theorem to derive